equipment costing $110000 with a salvage value of $15000 and an estimated life of 8 years has been

Kimmel, Accounting, 6e 134 Equipment costing $110000 with a salvage value of $15000 and an estimated life of 8 years has been depreciated using the straight-line rounde d) tor Years, Assurming a revised estimated total life of 6 years and no change in the salvage value, the depreciation expense $15833. $23750. O $11875. Open Show Work Click if you would like to Show Work for this question: LINK TO TEXT SAVE FOR LATER Question Attempts: 0 of 1 used 320 PM DOLL
Answer #1

Answer

  • All working forms part of the answer

A

Original Cost

$                                       110,000

B

Original Salvage value

$                                         15,000

C = A-B

Original Depreciable base

$                                         95,000

D

Original Life

8

E = C/D

Annual Depreciation as per SLM

$                                         11,875

F =E

Year 1 Depreciation expense

$                                         11,875

G = E

Year 2 Depreciation expense

$                                         11,875

H = A - F - G

Book Value at the time of revision of estimates

$                                         86,250

I

Salvage Value

$                                         15,000

J = H - I

Depreciable base [New]

$                                         71,250

K

New life [total]

6

L = 6 years - 2 years passed

Life remaining

4

K = J/L

Depreciation expense for Year 3

$                                         17,813

Correct Answer

Option #2: $ 17,813

  • Correct Answer Option #2: $ 17,813